OpenStudy (anonymous):

Inverse of (x+8)/(x)???

OpenStudy (anonymous):

no it doesnt

OpenStudy (anonymous):

they dont

OpenStudy (anonymous):

set x to y so x = (8+y)/y yx = 8+y = yx - y = 8 y(x-1)= 8 y = 8/(x-1) is your inverse

OpenStudy (anonymous):

thanks